OpenStudy (anonymous):

Given that the question 2x^2+x-4 has roots a and B, form a quadratic equation with integer coefficients and with roots (a-(1/2B)) and (B-(1/2a)) a+B= -1/2 aB=2

OpenStudy (anonymous):

i don't understand... what's the quesion?

OpenStudy (anonymous):

\[ a-\frac{1}{2}B=\frac{2a-B}{2}=r_1\\ B-\frac{1}{2}a=\frac{2B-a}{2}=r_2 \] \[r_1+r_2=\frac{a-B}{2}\\ r_1r_2=\frac{(2a-B)(2B-a)}{4}\] \[P(x)=(x-r_1)(x-r_2)=x^2-(r_1+r_2)x+r_1r_2\] \[ \begin{align} P'(x)&=x^2+\frac{B-a}{2}x+\frac{(2a-B)(2B-a)}{4}\\ &=4x^2+2(B-a)x+(2a-B)(2B-a) \end{align}\] This doesn't get anymore fun... Solve from here.

OpenStudy (anonymous):

From the original equation, we have that the sum of the roots is -1/2, and the product of the roots is -2, not positive 2.

OpenStudy (anonymous):

Hi the product of roots is positive 2 as i have accidentally wrote the wrong equation it is 2x^2+x+4:D Thank yiu

OpenStudy (anonymous):

May I ask why you removed my medal? It is not the goal of this site to give complete answers.